OpenStudy (anonymous):

Determine if (3/4,-2) is a solution to -4x-7y=11 . Please show all of your work

OpenStudy (anonymous):

-4(3/4) - 7(-2) = 11 (-12/4) - (-14) = 11 -3 - -14 =11 -3 + 14 = 11 11 = 11 (3/4,-2) IS a solution to -4x-7y=11